OpenStudy (jenniferkeryakous):

How do you solve: radical 49/8 < is in fraction form. How do you get the answer?

OpenStudy (babalooo):

\[\sqrt{\frac{ 49 }{ 8 }}\] is the same as \[\frac{ \sqrt{49} }{ \sqrt{8} } \]

OpenStudy (jenniferkeryakous):

@babaloo i know that part but what do i do next

OpenStudy (kkutie7):

\[\frac{ 7 }{ \sqrt{8} }\]

OpenStudy (kkutie7):

\[\sqrt{8}=2\sqrt{2}\]

OpenStudy (kkutie7):

\[\frac{ 7 }{ 2\sqrt{2} }\] yess?

OpenStudy (jenniferkeryakous):

@kkutie7 the calculator gave me a different answer: 7radical 2 /4 idk how it got it.

OpenStudy (aum):

Many books and teachers do not like radicals in the denominator and so you need to get rid of it by "rationalizing the denominator". Multiply top and bottom by sqrt(2): \[ \large \sqrt{\frac{49}{8}} = \frac{7}{\sqrt{4\times 2}} = \frac{7}{2\sqrt{2}} = \frac{7}{2\sqrt{2}} * \frac{\sqrt{2}}{\sqrt{2}} = \frac{7\sqrt{2}}{2*2} = \frac{7\sqrt{2}}{4} \]